# Elastic scattering at $\sqrt{s}=6$ GeV up to $\sqrt{s}=13$ TeV

### v1

@@ V.3 The real part of the elastic scattering amplitude
One of the origins of the non-linear behavior of the differential cross sections may arise the different $t$ dependence of the imaginary and real parts of the scattering amplitude. In most part, in the different approaches it is supposed that this $t$ dependence is the same for both parts of the scattering amplitude. It should be noted the importance of determining the size of the real part of the scattering amplitude was emphasized in many work of Andre Martin. If at the LHC the value of $\rho(s,t)$ is measured at high precision, it will give the possibility to check up the validity of the dispersion relations [79].
In the analysis of the experimental data [65] two cases were considered. One is so-called ”central” case in which the ratio of the real to imaginary parts of the scattering amplitude is independent of momentum transfer or slightly decreases. The other, so-called ”peripheral” case takes into account the assumption that $\rho(s,t)$ grows with momentum transfer. Really, the last case contradicts the dispersion relations; hence, it has non-physical motivation.
@@
For simplicity, very often they use, the so-called, local or derivative dispersion relations (see for example [45]) to determine the real part of the scattering amplitude. For example, the COMPETE Collaboration used
@@
In most cases, the real part is taken as proportional to the imaginary part of the scattering amplitude. So the slopes of both parts equals each other. There is also some unusual assumption about the growth of the real part of the scattering amplitude at small momentum transfer relative to its imaginary part (so-called - the peripheral case [65]). Obviously, both assumptions do not satisfy the dispersion relations, especially the last one.
.                 Refer to caption .                 Refer to caption
Figure 13: The ratio of the cross-odd part of the scattering amplitude of elastic proton-proton scattering to its cross even part (a) at 13 TeV; (b) at 9.26 GeV
.                 Refer to caption .                 Refer to caption
@@
Let us examine the origins of the complex $t$ dependence of the real part. Take the scattering amplitude in the form
$F(s,t)\ =\ hs^{\Delta}e^{BtLn(\hat{s})}.$ (43)
where the complex $\hat{s}$ is used only in the exponential. In this case, the real part will be negative (in Fig. 11, it is shown by the short dashed line). and essentially different from the behavior of the imaginary part (long-dashed line). Now let us use $\hat{s}$ is used in all parts of the scattering amplitude
$F(s,t)\ =\ h\hat{s}^{\Delta}e^{Bt\ln(\hat{s})}.$ (44)
The dispersion relations lead to the fact that the slope of the real part of the scattering amplitude must be larger than the slope of imaginary part. For example, if the imaginary part of the spin-nonflip hadron elastic scattering amplitude takes a simple exponential form $ImF_{+}\sim he^{Bt}$, then from eq.(V.3) we have that the real part of the $F_{+}$ will be $ReF_{+}=(1.+Bt)e^{Bt}$. Hence it has zero in the region of momentum transfer around 0.10.15-0.1--0.15- 0.1 - - 0.15 GeV2
At last, but not least, it should be noted that the unitarization procedure has a strong influence on the $t$ dependence of the real part of the scattering amplitude. For example, take the standard eikonal form of unitarization. If the Born term is taken in the ordinary exponential form (eq.44), then the imaginary part has a constant slope (see Fig. 12 [top]) and the real part has first zero at small momentum transfer. After eikonalization both slopes of the imaginary and real parts of the scattering amplitude have a strong $t$ dependence (Fig. 12 [bottom]).
VI New effects in diffraction elastic scattering at small angles
In the fitting procedure of the experimental data, only statistical errors were taken into account. As the systematic errors are mostly determined by indefiniteness of luminosity, they are taken into account as an additional normalization coefficient. This method essentially decreases the space of a possible form of the scattering amplitude. This allowed us to find the manifistation of some small effects at 13 TeV experimental data for the first time [19-21]. Our further researches with taking into account a wider range of experimental data confirm such new effects. We determined the new anomalous term with a large slope as
$f_{an}(t)=$ $ih_{an}\ln{(\hat{s}/s_{0})}/k$
$\exp[-\alpha_{an}(|t|+(2t)^{2}/t_{n})\ln{(\hat{s}/s_{0})}]\ F_{em}^{2}(t);$
where $h_{an}$ is the constant determining the size of the anomalous term with a large slope - $\alpha_{an}$; $F_{em}(t)$ is the electromagnetic form factor, which was determined from the GPDs [25], and $\ k=\ln(13000^{2}\ {\rm GeV}^{2}/s_{0})$ is introduced for normalization of $h_{an}$ at $13$ TeV, $t_{n}=1$ GeV2-normalization factor (see Appendix B for definitions of $\hat{s}$ and $s_{0}$). Such a form adds only two additional fitting parameters, and this term is supposed to grow with energy of order $\ln{(\hat{s}/s_{0})}$. The term has a large imaginary part and a small real part determined by the complex $\hat{s}$.
It is proportional to electromagnetics form factors, and the analysis of the experimental data above $6$ GeV gives the sizes of the constant $h_{an}=0.997\pm 0.006$. It is less than obtained in the fit of only high energy data, but now it has a more complicated energy dependence and a very small error.
Figure 15: The comparison of the HEGS model calculation of $d\sigma/dt$ of $pp$ scattering at $\sqrt{s}=27.4$ TeV [left]; and $p\bar{p}$ scattering at $\sqrt{s}=30$ GeV [right].
Figure 16: The comparison of the HEGS model calculation of $d\sigma/dt$ of $pp$ scattering at $\sqrt{s}=19.4$ TeV [left]; and at $\sqrt{s}=12.3$ GeV [right].
Figure 17: The comparison of the HEGS model calculation of $d\sigma/dt$ of $pp$ scattering at $\sqrt{s}=9.9$ TeV [left]; and at $\sqrt{s}=6.1$ GeV [right].
Our method helps us to find a small oscillation effect in the differential cross section at small momentum transfer. Such oscillation can be determined by an additional oscillation term in the scattering amplitude. Our fitting procedure takes the oscillatory function
$f_{osc}(t)=$ $ih_{osc}(1\pm i)(\ln{(\hat{s}/s_{0})}/k+h_{s}/\hat{s})\ J_{1}(\tau)/\tau\ A^{2}(t),$
$\tau=\pi\ (\phi_{0}-t)/t_{0};$
here $J_{1}(\tau)$ is the Bessel function of the first order; $t_{0}=1/[a_{p}/(\ln{(\hat{s}/s_{0})}/k)]$, where $a_{p}=17.15$ GeV-2 is the fitting parameter that leads to AKM scaling on $\ln{(\hat{s}/s_{0})}$; $A(t)$ is the gravitomagnetic form factor, which was determined from the GPDs [25], and $h_{osc}$ is the constant that determines the amplitude of the oscillatory term with the period determined by $\tau$. This form has only a few additional fitting parameters and allows one to represent a wide range of possible oscillation functions. The phase $\phi_{0}$ is obtained near zero and has a different sign for $pp$ and $p\bar{p}$ scattering. Inclusion in the fitting procedure of the data of $p\bar{p}$ elastic scattering shows that part of the oscillation function changes its sign for the crossing reactions. As a result, the plus sign is related with $pp$ and minus with $p\bar{p}$ elastic scattering. Hence, this part is the crossing-odd amplitude, which has the same simple form for $pp$ and $p\bar{p}$ scattering only with a different sign.
The wider energy region used in this analysis allows one to reveal the logarithmic energy dependence of the oscillation term. Let us compare the constant (size) of the oscillation function of three independent analyses (only $13$ TeV, all LHC data, all data above $500$ GeV). Now from our fitting procedure we obtain $h_{osc}=0.227\pm 0.007\ {\rm GeV}^{-2}$ The size of $h_{osc}$ is obtained as is smaller than obtained in the only high energy data; however, the error is decreased essentially. Perhaps, this reflects a more complicated form of energy dependence obtained now. Note, despite the logarithmic growth of the oscillation term, its relative contribution decreases as the main scattering amplitude grows as $\ln^{2}(s)$.
Figure 18: The HEGS model calculation of $d\sigma/dt$ of $pn$ scattering (top) at $\sqrt{s}=26$ GeV; (down) $\sqrt{s}=19.4$ GeV; [left] with TOTEM data [66] (requires the additional normalization of the data $k=0.91$); [right] with TOTEM data [96, 97] (with the additional normalization of the data $k=0.915$).
Figure 19: The HEGS model calculation of $d\sigma/dt$ of $pn$ scattering (top) at $\sqrt{s}=13.7$ GeV; (down) $\sqrt{s}=4.8$ GeV; [left] with TOTEM data [66] (requires the additional normalization of the data $k=0.91$); [right] with TOTEM data [96, 97] (with the additional normalization of the data $k=0.915$).
VII Proton-proton and proton-antiproton elastic scattering and Odderon contributions
At high energies and in the region of small momentum transfer the difference between the $pp$ and $p]bar{p}$ differential cross sections comes in most part from the CNI term, as the real part of the Coulomb amplitude has a different sign in these reactions. In the standard fitting procedure, one neglects the $\alpha^{2}$ term the equation takes the form:
$d\sigma/dt=$ $\pi[(F_{C}(t))^{2}\!+\!(\rho(s,t)^{2}+1)(ImF_{N}(s,t))^{2})$ (47)
$+$ $2(\rho(s,t)+\alpha\varphi(t))F_{C}(t)ImF_{N}(s,t)],$
[…]
It is supposed that the real part of the $pp$ and $p\bar{p}$ scattering have the same sign, at least at small $t$. Such a difference, obtained in the HEGS model is presented in Fig. 6, where the comparison of the HEGS model calculation of $d\sigma/dt$ of $pp$ scattering at $\sqrt{s}=30.6$ GeV (solid line and full squares) and $\bar{p}p$ at $\sqrt{s}=30.4$ GeV (dashed line and open triangles) is shown. In both cases the additional normalization $n=1$.
@@ VII Proton-proton and proton-antiproton elastic scattering and Odderon contributions
Of course, the Odderon amplitude changes the size and $t$ dependence of the real par of the full amplitude. In the recent paper [91] the different cases (with and without Odderon contributions) were analysed. It was noted that the effect of incorporating the Odderon becomes notably significant when analysing specific subsets of data. It is remarkable, that the authors note ”we will get too large $\rho^{\bar{P}p}$ at $\sqrt{s}\sim 541$ GeV in disagreement with the data UA4/2. As a result, they restore an old problem of the value of $\rho(s,t)$ at $\sqrt{s}\sim 540$ GeV. However, as was noted in the introduction, this problem strongly depends on the form of the non-linear slope and can be solved in [73]. Now, in the present model, the Odderon amplitude essentially decreases as $t\rightarrow 0$. The value of $\rho(\sqrt{s}=541\rm{GeV},t=0)=0.122$, which coincides with the result of the UA4/2 Collaboration.
@@
In Fig. 13, the ratio of the cross-odd part of the scattering amplitude of elastic proton-proton scattering to its cross even part at 13 TeV and at 9.26 GeV is presented. It is seen that this ratio is small at 13 TeV, except for the position of the diffraction minimum. At low energy, this ratio is small at small t but increases essentially in the region of large momentum transfer.
@@
However, the Odderon contribution is very important at the diffraction minimum. The real part of the scattering amplitude, which is positive at $t=0$ and $\sqrt{s}>30$ GeV, changes its sign at larger $t$ thus corresponding to the dispersion relations, but, in any case it fulfills the diffraction dip in the differential cross sections. In Fig. 14, it is shown the sizes of the differential cross sections in the region of the diffraction minimum with and without the Odderon contributions at $\sqrt{s}$ around $53$ GeV and at $10$ TeV.

### v3

@@ V.3 The real part of the elastic scattering amplitude
It is interesting that the form of the real part of the hadron elastic scattering amplitude is similar to its imaginary part. Of course, they are not proportional to each other as their connection has to satisfy the dispersion relations [30] which require, for example, the changing size of the real part at sufficiently small momentum transfer. In fact, in Fig. 14, we see that the real part changes its sign essentially earlier than the position of the diffraction minimum. At LHC energies this happens in the area of momentum transfer approximately equal to $0.2$ GeV2. But before that, we can see the crossing point at $|t|\sim 0.06$ GeV2. Likely, the behavior of the imaginary part the slope of the real part at very small momentum transfer is also practically proportional to the size of the total cross section at different energies and grows at larger momentum transfers. It can be see that the real part at LHC energies has the negative maximum at approximately $|t|=0.3$ GeV2 situated near the diffraction minimum. Hence, it essentially impacts the form and size of the diffraction minimum in the differential cross sections.
Especially the size and energy dependence of the real part of the hadronic amplitude impact the differential cross sections in the CNI region. The model descriptions of the differential cross section at a small momentum transfer and at low energies are presented in Figures 5, 15-17. These figures show the experimental data with high precision and only statistical errors, which were taken into account in out fitting procedure. Obviously, the model reproduces the experimental data very well in a wide energy region.
One of the origins of the nonlinear behavior of the differential cross sections may be the different $t$ dependencies of the imaginary and real parts of the scattering amplitude. In most part, in different approaches it is supposed that this $t$ dependence is the same for both parts of the scattering amplitude. It should be noted the importance of determining the size of the real part of the scattering amplitude was emphasized in many work of Andre Martin. If at the LHC the value of $\rho(s,t)$ is measured at high precision, it gives the possibility to check up the validity of the dispersion relations [56].
In the analysis of the experimental data [93] two cases were considered. One is the so-called ”central” case, in which the ratio of the real to imaginary parts of the scattering amplitude is independent of momentum transfer or slightly decreases. The other, the so-called ”peripheral” case, takes into account the assumption that $\rho(s,t)$ grows with momentum transfer. Really, the last case contradicts the dispersion relations; hence, it has non-physical motivation.
@@
For simplicity, very often they use the so-called local or derivative dispersion relations (see, for example, [45]) to determine the real part of the scattering amplitude. For example, the COMPETE Collaboration used
@@
Figure 16: The size of the $\rho(s,t)$ ratio of the real to imaginary part of the $pp$ scattering amplitude is calculated in the model at the energy $\sqrt{s}=0.51,7,13$ TeV (solid, short dashed and dashed line respectively) depending on $s$ and $t$
In most cases, one assumes that the real part is taken proportional to the imaginary part of the scattering amplitude. So the slopes of both parts are equal to each other. There is also some unusual assumption about the growth of the real part of the scattering amplitude at a small momentum transfer relative to its imaginary part (the so-called the peripheral case [93]). Obviously, both assumptions do not satisfy the dispersion relations, especially the last one.
Let us examine the origins of the complicated $t$ dependence of the real part. Take the scattering amplitude in the form
$F(s,t)\ =\ hs^{\Delta}e^{BtLn(\hat{s})}.$ (41)
where the complex $\hat{s}$ is used only in the exponential. In this case, the real part is negative (in Fig. 13, it is shown by the short dashed line) and essentially differs from the behavior of the imaginary part (long-dashed line). Now let us use $\hat{s}$ in all parts of the scattering amplitude
$F(s,t)\ =\ h\hat{s}^{\Delta}e^{Bt\ln(\hat{s})}.$ (42)
The dispersion relations lead to the fact that the slope of the real part of the scattering amplitude must be larger than the slope of the imaginary part. For example, if the imaginary part of the spin-nonflip hadron elastic scattering amplitude takes a simple exponential form $ImF_{+}\sim he^{Bt}$, then from eq.(40) we have that the real part of the $F_{+}$ is $ReF_{+}=(1.+Bt)e^{Bt}$. Hence, it has zero in the region of momentum transfer around 0.10.15-0.1--0.15- 0.1 - - 0.15 GeV2
At last, but not least, it should be noted that the unitarization procedure has a strong influence on the $t$ dependence of the real part of the scattering amplitude. For example, take the standard eikonal form of unitarization. If the Born term is taken in the ordinary exponential form (eq.42), then the imaginary part has a constant slope (see Fig. 15 [top]) and the real part has the first zero at a small momentum transfer. After eikonalization both slopes of the imaginary and real parts of the scattering amplitude have a strong $t$ dependence (Fig. 15) [bottom]).
VI The value of $\rho(s,t)$ and Odderon contributions
In Fig. 16, the ratio $\rho(s,t)$ of the real to imaginary part of the hadron elastic scattering amplitude is presented for different energies. Such a complicated structure of $\rho(s,t)$ is determined by the changes of the sign of the real and imaginary parts of the scattering amplitude. At a small momentum transfer, the size of $\rho(s,t)$ is small as the real part changes its sign. Contrary, when the imaginary part changes its sign, the size of $\rho(s,t)$ grows very faster. The energy dependence of $\rho(s,t)$ is due to the movement of the position of the diffraction minimum, and hence to the energy dependence of the imaginary part of the scattering amplitude.
At high energies and in the region of small momentum transfer the difference between the $pp$ and $p\bar{p}$ differential cross sections comes in most part from the CNI term, as the real part of the Coulomb amplitude has a different sign in these reactions. In the standard fitting procedure, one neglects the $\alpha_{em}^{2}$ term and then the equation takes the form:
$d\sigma/dt=$ $\pi[(F_{C}(t))^{2}\!+\!(\rho(s,t)^{2}+1)(ImF_{N}(s,t))^{2})$
$+$ $2(\rho(s,t)+\alpha\varphi(t))F_{C}(t)ImF_{N}(s,t)],$ (43)
[…]
Figure 17: $\rho(s,t=0)$: dashed line - $p\bar{p}$ scattering, solid line -$pp$ scattering; the same but without the second Rejions contributions (tiny dashed - $p\bar{p}$ and long dashed $pp$ scattering).
Figure 18: The ratio $R(s,t)$ of the cross-odd part of the scattering amplitude of elastic proton-proton scattering to its cross even part at 13 TeV (dashed line) and at 9.26 GeV (solid line).
@@ VI The value of $\rho(s,t)$ and Odderon contributions
The Odderon amplitude has a large real part compared to it imaginary part. Hence, the odderon contribution changes the size and $t$ dependence of the real part of the full amplitude. In a recent paper [94] different cases (with and without Odderon contributions) were analysed. It was noted that the effect of incorporating the Odderon becomes notably significant when analysing specific subsets of data. It is remarkable that the authors note ”we will get too large $\rho^{p\bar{p}}$ at $\sqrt{s}\sim 541$  GeV in disagreement with the data UA4/2. As a result, they restored an old problem of the value of $\rho(s,t)$ at $\sqrt{s}\sim 540$ GeV. However, as was noted in the introduction, this problem strongly depends on the form of the non-linear slope and can be solved in [95, 29]. Now, in the present model, the Odderon amplitude essentially decreases as $t\rightarrow 0$. The value of $\rho(\sqrt{s}=541\rm{GeV},t=0)=0.122$, which coincides with the result of the UA4/2 Collaboration.
Figure 20: The contributions $\Delta(d\sigma/dt)$ of the anomalous term with a large slope to the differential cross sections of $pp$ scattering at $\sqrt{s}=19$ GeV (solid line) and at $\sqrt{s}=13$ TeV.
@@
In Fig. 18, the ratio of the cross-odd part of the scattering amplitude in the HEGS model of elastic proton-proton scattering to its cross even part at 13 TeV and at 9.26 GeV is presented. It is seen that this ratio is small at 13 TeV, except for the position of the diffraction minimum. At low energy, this ratio is small at small t but increases essentially in the region of a large momentum transfer.
@@
However, the Odderon contribution is very important at the diffraction minimum. The real part of the scattering amplitude, which is positive at $t=0$ and $\sqrt{s}>30$ GeV, changes its sign at larger $t$, thus corresponding to the dispersion relations but, in any case, it fulfills the diffraction dip in the differential cross sections. Figure 17 shows the sizes of the differential cross sections in the region of the diffraction minimum with and without the Odderon contributions at $\sqrt{s}$ around $53$ GeV and at $10$ TeV.
VII New effects in diffraction elastic scattering at small angles
In the fitting procedure of the experimental data, only statistical errors were taken into account. As the systematic errors are mostly determined by indefiniteness of luminosity, they are taken into account as an additional normalization coefficient. This method essentially decreases the space of a possible form of the scattering amplitude. This allowed us to find the manifestation of small effects at 13 TeV experimental data for the first time [19-21]. Our further researches with taking into account a wider range of experimental data confirm the existence of such new effects. We determined the new anomalous term with a large slope as
$f_{an}(t)=$ $ih_{0an}\ln{(\hat{s}/s_{0})}/k(1+h_{1an}\ln{(\hat{s}/s_{0})}/k)$ (44)
$\exp[-\alpha_{an}(|t|+(2t)^{2}/t_{n})\ln{(\hat{s}/s_{0})}]\ G^{2}_{em}(t);$
where $h_{an}$ is the constant determining the size of the anomalous term with a large slope - $\alpha_{an}$; $G^{2}_{em}(t)$ is the electromagnetic form factor, which was determined from the GPDs [76], and $\ k=\ln(13000^{2}\ {\rm GeV}^{2}/s_{0})$ is introduced for normalization of $h_{an}$ at $13$ TeV, and $t_{n}=1$ GeV2 is the normalization factor. This form adds only two additional fitting parameters, and this term is supposed to grow with energy of order $\ln{(\hat{s}/s_{0})}$. The analysis of the data in a wide energy region gives a complicated energy dependence which is reflected in eq.(44) The term has a large imaginary part and a small real part determined by the complex $\hat{s}$. It is remarkable that the anomalous slope is determined with high accuracy $\alpha_{an}=0.513\pm 0.004$ GeV-2 from the analysis of the whole sets of experimental data.
Figure 21: The absolute value of the oscillation term $(Re^{2}F_{os}+Im^{2}F_{os})^{1/2}$ at $\sqrt{s}=13$ TeV (solid line $pp$ scattering; dashed points line - $p\bar{p}$ scattering ). and at $\sqrt{s}=540$ GeV (dotted line $pp$ scattering; dashed line - $p\bar{p}$ scattering)
Figure 22: The HEGS model calculation of $d\sigma/dt$ of $pn$ scattering
at $\sqrt{s}=26$ GeV; [right] $\sqrt{s}=19.4$ GeV.
It is proportional to electromagnetic form factors, and the analysis of the experimental data above $6$ GeV gives the size of the constant $h_{an}=1.13\pm 0.04$. It is a little less than obtained in the fit of only high energy data but now it has a more complicated energy dependence and a very small error. It impacts the differential cross sections at small $t$ and the size of $\sigma_{tot}(s)$. However, it practically does not influence the value of $\rho(s,t)$ as the imaginary and real parts of the scattering amplitude grow proportionally. The contributions of the anomalous term with a large slope to the differential cross sections of $pp$ scattering -
$\Delta(d\sigma/dt=d\sigma/dt|_{with}-d\sigma/dt|_{without})$
are shown in Fig. 20 at $\sqrt{s}=19$ GeV (solid line) and at $\sqrt{s}=13$ TeV (dashed line). It can be seen that the contribution of the anomalous term disappears after $-t>0.2$ GeV2. At low energies its relative contribution is less but remains visible and decreases with $t$ slowly.
Our method helps us to find a small oscillation effect in the differential cross section at small momentum transfer [31]. Such oscillation can be determined by an additional oscillation term in the scattering amplitude [96]. The theoretical predictions of some oscillations in the differential cross section has a long story, starting from one of the first paper [97] taking into account a multi pomeron exchange and the papers basing on the violation of the Pomeranchuk theorem (for example, [6]) up to a recent paper [98]. In our fitting procedure the oscillatory function is
$f_{osc}(t)=$ $ih_{0osc}(1\pm ih_{1osc})$
$(\ln{(\hat{s}/s_{0})}/k+h_{s}/\hat{s})\ J_{1}(\tau)/\tau\ A^{2}(t),$
$\tau=\pi\ (\phi_{0}-t)/t^{\prime}_{0};$
here $J_{1}(\tau)$ is the Bessel function of the first order; $t_{0}=1/[a_{p}/(1+2\ln^{2}{(\hat{s}/s_{0})}/k)]$, where $a_{p}=17.15$ GeV-2 is the fitting parameter that leads to the AKM scaling on $\ln{(\hat{s}/s_{0})}$; $A(t)$ is the gravitomagnetic form factor, which was determined from the GPDs [76], and $h_{osc}$ is the constant that determines the amplitude of the oscillatory term with the period determined by $\tau$. This form has only a few additional fitting parameters and allows one to represent a wide range of possible oscillation functions. The phase $\phi_{0}$ is obtained near zero for $pp$ scattering and it is slightly small and negative for $p\bar{p}$ scattering, so it has a different sign for $pp$ and $p\bar{p}$ scattering. The inclusion of the $p\bar{p}$ elastic scattering data in the fitting procedure shows that part of the oscillation function changes its sign for the crossing reactions. As a result, the plus sign is related to $pp$ and minus with $p\bar{p}$ elastic scattering. Hence, this part is the crossing-odd amplitude, which has the same simple form for $pp$ and $p\bar{p}$ scattering only with a different sign.
Figure 23: The HEGS model calculation of $d\sigma/dt$ of $pn$ scattering
at $\sqrt{s}=13.7$ GeV; [right] $\sqrt{s}=4.8$ GeV.
The wider energy region used in this analysis allows one to reveal the logarithmic energy dependence of the oscillation term. The constant (size) of the oscillation function was determined from experimental data at high energies (all data above $500$ GeV) as $h_{osc}^{c}=0.270\pm 0.007\ {\rm GeV}^{-2}$, whereas in our analysis of the whole sets of the data $\sqrt{s}>3.6$ GeV. the $h_{osc}^{c}=0.232\pm 0.009\ {\rm GeV}^{-2}$ The size of $h_{osc}$ is obtained as is smaller than obtained in the only high energy data; however, the error is decreased essentially. Perhaps, this reflects a more complicated form of the energy dependence obtained now. Note, despite the logarithmic growth of the oscillation term, its relative contribution decreases as the main scattering amplitude grows as $\ln^{2}(s)$.
The oscillation term is represented in Fig. 21 at different energies $\sqrt{s}=13$ TeV and $\sqrt{s}=540$ TeV for the proton-proton and proton-antiproton scattering. It can be seen that the amplitude of the oscillation term increases at high energies, but the period of oscillations decreases with growing energy. Especially note that the period of oscillation is essentially less for $p\bar{p}$ than for $pp$ scattering. This can explain the difference which was noted in [99], in the behaviour of oscillations of the differential cross sections at $\sqrt{s}=540$ GeV, which were found in [100] and at $\sqrt{s}=13$ TeV [31]. Of course, the amplitude of oscillations is weak, so it is very difficult to identify it by experiment. However, our analysis of a wide energy region of experimental data gives a small error in the definition of the parameter determining the size of oscillations. This shows that further meticulous work is required, both from the theoretical and experimental points of view.